OpenStudy (anonymous):

each interior angle of a regular polygon is twice the measure of each exterior angle. How many diagonals does the polygon have?

jimthompson5910 (jim_thompson5910):

are you able to figure out how many sides there are?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

If x is the exterior angle and y is the interior angle then x+y = 180 "each interior angle of a regular polygon is twice the measure of each exterior angle", so y = 2x

jimthompson5910 (jim_thompson5910):

x+y = 180 x + 2x = 180 x = ???

OpenStudy (anonymous):

180?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

3x = 180 x = ??

OpenStudy (anonymous):

360?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

3x means 3 times x

jimthompson5910 (jim_thompson5910):

3 times _____ = 180 what goes in the blank?

OpenStudy (anonymous):

60?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

x = 60 2x = ??

OpenStudy (anonymous):

120?

jimthompson5910 (jim_thompson5910):

so each interior angle is 120 degrees each exterior angle is 60 degrees

jimthompson5910 (jim_thompson5910):

if each exterior angle is 60 degrees, then how many sides are there?

OpenStudy (anonymous):

120?

jimthompson5910 (jim_thompson5910):

120 sides?

OpenStudy (anonymous):

idont get it? ill have to divide? 60?

jimthompson5910 (jim_thompson5910):

n = 360/E n = 360/60 n = ???

OpenStudy (anonymous):

aah! 6 ? :>

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now use the formula d = n*(n-3)/2 d = # of diagonals n = number of sides

OpenStudy (anonymous):

9?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

yeay so it has 9 diagonals? right? :>

jimthompson5910 (jim_thompson5910):

yes it does

OpenStudy (anonymous):

yeaay! thank you very much :) ur soo smart :) i wish i was like you. Lol thank you bro :)

jimthompson5910 (jim_thompson5910):

you're welcome